OpenStudy (anonymous):

hello every one can you solve this equation 11/4 over m+4=7 over m+4

OpenStudy (anonymous):

is your problem this: (11/4)/m + 4 = 7 / (m+4)

OpenStudy (anonymous):

this is the equation 1-1/4 over m+4=7/m+4

OpenStudy (anonymous):

not 11

OpenStudy (anonymous):

Can you put your equation with parenthesis such that I can better understand your order of operations?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[1\frac{ 1 }{ 4 }=\frac{ 7 }{ m+4 }\]

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

thats wrong \[1-\frac{ 1 }{ 4 }=\frac{ 7 }{ m+4 }\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes,all equation is over m+4

OpenStudy (anonymous):

alright then... lets begin by simplifying the right side of the equation. What is 1 - (1/4) ?

OpenStudy (anonymous):

hello kwo8 those first equation is over m+4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

therefore, your problem is: ((1-(1/4))/(m+4)) = 7/(m+4)

OpenStudy (anonymous):

I hope I am getting this problem correct because I can't start until I'm sure i've got the problem right.

OpenStudy (anonymous):

kwo8 you asked me what is 1(1/4) i just know is 4/4 because i multiply 4 into 1 angd divide 4 to one

OpenStudy (anonymous):

lets go back to the first step. I want to reconfirm that your original problem is ((1-(1/4))/(m+4)) = 7/(m+4)

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

I would like to comment that this problem is very strange. That is because there is m+4 on both sides of the equation. This means that you can cancel it out... rendering this problem meaningless to solve. Are you sure this is right!?

OpenStudy (anonymous):

sorry kwo8 if i'm interupting you and increase the problem

OpenStudy (anonymous):

yes i'm sure

hartnn (hartnn):

it might be \(\huge1-\frac{1/4}{m+4}=\frac{7}{m+4}\)

OpenStudy (anonymous):

good point... is hartnn correct?

hartnn (hartnn):

if not please take the snapshot of question and post

OpenStudy (anonymous):

kwo can you see those equation by hartnn that correct

OpenStudy (anonymous):

hello hartnn those equation is correct

OpenStudy (anonymous):

ok... we can start solve the problem now that we have it. First you should move all numbers related to 'm' to one side of the equation. This means that you will have to move (1/4)/(m+4) to the right side of the equation. What operation are you doing now? (subtraction? addition? division? multiplication?)

OpenStudy (anonymous):

multiplycation

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):